C14 #2 Weygandt, Accounting Principles, 13e Help System Announcements Exercise 14-19 a-b At December 31, Pharoah Corporation can you work it out for me also
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Pharaoh Corporation
Answer to a
Calculation of Book Value per share
Amount $
Total Equity    3,100,000.00
Less:
Preferred Stock value       750,000.00
Value of Common stock and retained earnings 2,350,000.00 N
Number of Common Shares outstanding       250,000.00 L
Book Value per share                   9.40 O=N/L
So book value per share is $ 9.40.
Answer to b
Calculation of arrear Preferred dividend Note
Preferred Stock value       750,000.00 A
Dividend rate 7% B
Arrear Dividend         52,500.00 C=A*B
Calculation of excess payment in case of redemption of Preferred Shares
Call value per share                 60.00 D
Par Value                 50.00 E
Difference                 10.00 F=D-E
Number of Preferred Shares         15,000.00 G
Total value       150,000.00 H=F*G
Less: Paid in capital in excess of Par Preferred Stock         49,800.00 I
Excess payment upon redemption of Preferred Shares       100,200.00 J=H-I
Calculation of Book Value per share
Amount $
Total Equity    3,100,000.00
Less:
Preferred Stock value       750,000.00
Paid in capital in excess of par Preferred Stock         49,800.00
Arrear Dividend         52,500.00 See C
Excess payment upon redemption of Preferred Shares       100,200.00 See J
Value of Common stock and retained earnings 2,147,500.00 K
Number of Common Shares outstanding       250,000.00 See L
Book Value per share                   8.59 M=K/L
So book value per share is $ 8.59.